Solve the formula for the given variable. A = bh; h (Geometry)
Answer #1

Given that

We need to solve the formula for the given equation

Given equation is A=bh

It is an literal equation we need to solve to find "h".

A=bh

Move " b" to right side

I.e. h=A/b

There h=A/b

It is the solution for literal equation A=bh

"""" h=A/b"""""""